OpenStudy (anonymous):

What is the Molarity (M) of the HCl solution if 250 mL of the HCl solution reacts with magnesium to produce 4.20 L of H2 gas measured at 760 mmHg and 0 Celsius? Use the following unbalanced equation: Mg (s) + HCl (aq) --> MgCl2 (aq) + H2 (g) I balanced the equation to: Mg (s) + 2HCl --> MgCl2 (aq) + H2 (g) I think I start with the 4.20 L H2 and convert using (STP) 22.4 L H2= 1 mol H2... and then I'm stuck.

OpenStudy (anonymous):

ok so you start with: n= V/Vm and you get that n(H2) = 0,1875 mol now you get the ratio from the equation n(HCl)/n(H2) = 2/1 = 2 so n(HCl) = 2 * n(H2) = 0,375 mol now you just use: c=n/V c(HCl) = 0,375 mol / 250 * 10^-3 dm^3 = 1,5 M

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

not a problem :)